# High-order van Hove singularities and nematic instability in a kagome superconductor CsTi3Bi5

### v1

@@ III.4 Filling dependent electron-phonon interaction and superconductivity
The Eliashberg spectral function in Fig. 4(a) reveals four prominent peaks: one in the low-frequency region and the other three in the high-frequency region. This gives a cumulative electron-phonon coupling strength $\lambda=0.34$, indicating a low electron-phonon coupling regime of CsTi3Bi5. While all phonon bands contribute to the total electron-phonon coupling, the contribution of each mode varies significantly. The low-frequency modes account for $\sim 68$% of the total electron-phonon coupling (Fig 4(a)). The distribution of $\lambda_{\nu{\mathbf{q}}}$ in the $k_{z}=0$ plane is shown in Fig 4(b), which reveals substantial variation of $\lambda_{\nu{\mathbf{q}}}$ in the entire plane. The associated superconducting critical temperature $T_{c}$, calculated using the modified McMillan formula [53],

### v3

@@ III.4 Filling dependent electron-phonon interaction and superconductivity
The Eliashberg spectral function in Fig. 4(a) reveals four prominent peaks: one in the low-frequency region and the other three in the high-frequency region. This gives a cumulative electron-phonon coupling strength $\lambda=0.33$, indicating a low electron-phonon coupling regime of CsTi3Bi5. While all phonon bands contribute to the total electron-phonon coupling, the contribution of each mode varies significantly. The low-frequency modes account for $\sim 68$% of the total electron-phonon coupling (Fig 4(a)). The distribution of $\lambda_{\nu{\mathbf{q}}}$ in the $k_{z}=0$ plane is shown in Fig 4(b), which reveals substantial variation of $\lambda_{\nu{\mathbf{q}}}$ in the entire plane. The associated superconducting critical temperature $T_{c}$, calculated using the modified McMillan formula [69],